What is a Capital Asset?, Holding Period, Calculation of Gain or Loss, and Net Capital Gains (LO 4.1, 4.2, 4.3) During 2018, Tom sold Sears stock for $20,200. The stock was purchased 4 years ago for $28,280. Tom also sold Ford Motor Company bonds for $70,700. The bonds were purchased 2 months ago for $60,095. Home Depot stock, purchased 2 years ago for $2,020, was sold by Tom for $3,030.

-Calculate Tom's net gain or loss, and indicate the nature of the gain or loss.

2.Calculation of Gain or Loss and Net Capital Losses (LO 4.5)

In 2018, Michael has net short-term capital losses of $1,500, a net long-term capital loss of $27,000, and other ordinary taxable income of $45,000.

If an amount is zero, enter "0".

a. Calculate the amount of Michael's deduction for capital losses on his tax return for 2018.

b.b. Calculate the amount and nature of his capital loss carryforward.
$ Long-term capital loss carryforward
$ Short-term capital loss carryforward
Answer #1

Capital assets are those assets which has a useful life longer than 1 year and which is not intended to sale in regular course of business operation.For example a computer bought by the company to use it in the office is a capital asset for the company.

Holding period is the period for which the company has hold the asset i.e. Period between purchase date and sale date. An asset used for more than 1 year is called as long term asset whereas the assets used for less than 1 year are called as short term assets.

Part 1

In the given case.

Tom net gain loss calculation -

Sears stock purchased 4 years ago hence it would be of long term asset

Net long term loss = 20,200 - 28,280 = (8.080)

Ford Motor Company Bonds purchased 2 months ago would be short term asset

Net short term gain = 70,700 - 60,095 = 10,605

Home Depot Stock purchased 2 years ago would be long term asset

Net Long term gain = 3,030 - 2,020 = 1,010

Therefore total net long term gain will be 10,605 - 8,080 = 2,525

Total net short term gain will be 1,010.

Part 2

A) As per the Scehdule D of form 1040 tax return capital losses can be deducted to the with the capital gains.But if you have more capital losses than gain then we can caim upto only $ 3,000 against ordinary income using the short term losses first.

So in the given case there is no capital gain so the maximum amount Michael can deduct his losses from ordinary income is $ 3,000 in his tax return 2018.

B) As per the Schedule capital loss can be carry over indefinitely till it become zero by claiming $3,000 every year where no capital gain has arise.Therefore Micheal can carryforward the loss indefinitely

Also while offsetting the capital losses against ordinary income short term losses are used first.Therefore short term capital loss of 1500 will be fully used against ordinary income and other 1,500 will be used from long term capital loss.

Long Term capital loss to be carryforward- 27,000 -1,500 = $25,500

Short term capital loss to be carryforward - 1,500 - 1,500 = 0